# MetaGFN: Exploring Distant Modes with Adapted Metadynamics for Continuous GFlowNets

### v1

@@ Appendix G Experiment details: alanine dipeptide environment
Computing the Free Energy Surface To obtain a ground-truth free energy surface (FES) in $\phi$-$\psi$ space, we ran a 250ns NPT well-tempered metadynamics MD simulation of alanine dipeptide at temperature 300K ($\beta=0.4009$), pressure 1bar with the TIP3P explicit water model Jorgensen et al. [1983]. We used the PLUMMED plugin Bonomi et al. [2019] for OpenMM Eastman et al. [2017] with the AMBER14 force field Salomon-Ferrer et al. [2013].
@@
Parameterisation We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and $\hat{f}$ through three heads of an MLP with 3 hidden layers with 512 hidden units per layer, with GeLU activations and dropout probability $0.2$, similar to the Line Environment. The first two heads have output dimensions of $15$, parameterising the $6$ means, $6$ concentrations, and $3$ weights of the mixture of von Mises policy. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The means are mapped to the range $(-\pi,\pi)$ through $2\arctan(\cdot)$. Concentrations are parameterised in log space and passed through a sigmoid to map to the range $\ln(\kappa)\in(0,5)$. Mixture weights are normalised with the softmax function.
Replay buffer The replay buffer has capacity for $10^{4}$ trajectories. Trajectories are stored in the replay buffer only if the terminal state’s reward exceeds $10^{-10}$. When drawing a replay buffer batch, trajectories are bias-sampled: $50\%$ randomly drawn from the upper $30\%$ of trajectories with the highest rewards, and the remaining $50\%$ randomly drawn from the lower $70\%$.
Noisy exploration We use the same noise profile as the Line Environment, equation (13). The noise $\bar{\sigma}$ is now added to the concentration parameter $\kappa$. Concentration is related to standard deviation through $\sigma=\frac{1}{\kappa^{2}}$.
Thompson Sampling We use 10 heads with the bootstrap probability parameter set to 0.3.
@@
MetaGFN We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.01$, $n=2$, $\beta=0.4009$, $\gamma=0.1$, $w=10^{-5}$, $\kappa=10$, $\epsilon=10^{-6}$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.1$. Initial samples are drawn from a Gaussian distribution, mean centred $P_{||}$, variance $\sigma^{2}=(0.1,0.1)$, and initial momenta from a Gaussian, mean $\mu=(0,0)$, variance $\sigma^{2}=(0.05,0.05)$.
Training parameters The same as for the Line Environment, see Appendix E.1.
@@
Evaluation The L1 error of a histogram of on-policy samples, $\hat{\rho}(\phi,\psi)$, is computed via a two-dimensional generalisation of (14); $\text{error}=\frac{1}{2}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left|\hat{\rho}(\phi,\psi)-\frac{r(\phi,\psi)}{Z}\right|\mathrm{d}\phi\mathrm{d}\psi$, estimated by a discrete sum with grid spacing $0.1$.
Compute resources Experiments are performed in PyTorch on a desktop with 32Gb of RAM and a 12-core 2.60GHz i7-10750H CPU. It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches. The additional computational expense of running Adaptive Metadynamics was less than $10\%$.

### v2

@@ D.3 Alanine dipeptide environment details
Computing the free energy surface: To obtain a ground-truth free energy surface (FES) in $\phi$-$\psi$ space, we ran a 250ns NPT well-tempered metadynamics MD simulation of alanine dipeptide at temperature 300K ($\beta=0.4009$), pressure 1bar with the TIP3P explicit water model [47]. We used the PLUMMED plugin [48] for OpenMM [49] with the AMBER14 force field [50].
@@
Parameterisation: The forward and backward policies are a mixture of three bivariate von Mises distributions. We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and $\hat{f}$ through three heads of an MLP with 3 hidden layers with 512 hidden units per layer, with GeLU activations and dropout probability $0.2$. The first two heads have output dimensions of $15$, parameterising the $6$ means, $6$ concentrations, and $3$ weights of the mixture of von Mises policy. The third head has output dimension $1$ and parameterises the flow function $\hat{f}$. The means are mapped to the range $(-\pi,\pi)$ through $2\arctan(\cdot)$. Concentrations are parameterised in log space and passed through a sigmoid to map to the range $\ln(\kappa)\in(0,5)$. Mixture weights are normalised with the softmax function. We use batch size $b=64$. It takes approximately $10$ hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
@@
MetaGFN: We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.01$, $n=2$, $\beta=0.4009$, $\gamma=0.1$, $w=10^{-5}$, $\kappa=10$, $\epsilon=10^{-6}$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.1$. Initial samples are drawn from a Gaussian distribution, mean centred $P_{||}$, variance $\sigma^{2}=(0.1,0.1)$, and initial momenta from a Gaussian, mean $\mu=(0,0)$, variance $\sigma^{2}=(0.05,0.05)$. In Figure 12 we show the resulting learnt KDE potential with these parameters.
@@
Evaluation: The L1 error of a histogram of on-policy samples, $\hat{\rho}(\phi,\psi)$, is computed via a two-dimensional generalisation of (D.2); $\text{error}=\frac{1}{2}\int_{-\pi}^{\pi}\int_{-\pi}^{\pi}\left|\hat{\rho}(\phi,\psi)-\frac{r(\phi,\psi)}{Z}\right|\mathrm{d}\phi\mathrm{d}\psi$, estimated by a discrete sum with grid spacing $0.1$.
D.4 Grid environments details
Parameterisation: For the two-dimensional grid, we use a mixture of $4$ Gaussians. For the three and four dimensional (hyper)grids, we use a mixture of $2$ Gaussians (but use longer trajectories, see Section 4). We parameterise $\hat{p}_{F}$, $\hat{p}_{B}$ and the flow $\hat{f}$ through an MLP with 3 hidden layers, 512 hidden units per layer, GELU activation function and dropout probability $0.2$. The means and standard deviations are mapped to the range $\mu\in(-15,15)$ and $\sigma\in(0.1,7)$ through sigmoid functions. Mixture weights are normalised with the softmax function. We use batch size $b=128$. It takes approximately 10 hours to train a continuous GFlowNet in this environment with $B=10^{5}$ batches.
MetaGFN: We use $\texttt{freqRB}=2$, $\texttt{freqMD}=10$, $\Delta t=0.35$, $n=3$, $\beta=1$, $\gamma=2$, $w=0.10$, $\sigma=2$. The bias and KDE potentials are stored on a uniform grid with grid spacing $0.075/0.3/0.6$ (2, 3, and 4 dimensions respectively). Initial samples are drawn from a Gaussian distribution, mean centred at the origin, isotropic variance $\sigma^{2}=1$. Momenta are initialised to zero.
Evaluation: The L1 error between the histogram of the terminal states of $10^{4}$ on-policy samples and the exact distribution is computed, e.g. for the 2D grid, $\text{error}=\frac{1}{2}\int_{-15}^{15}\int_{-15}^{15}\left|\hat{\rho}(x,y)-\frac{r(x,y)}{Z}\right|\mathrm{d}x\mathrm{d}y$, estimated by a discrete sum with grid spacing $0.075$.